A firm in a competitive market has a cost function: TC = 86 + 12Q - 1.6Q2 + 0.1Q3. What is the firm's shutdown price.
Answer #1

Cost function is TC = 86 + 12Q - 1.6Q2 + 0.1Q3

Variable cost is VC = 12Q - 1.6Q2 + 0.1Q3

Average Variable cost is AVC = VC/Q = 12 - 1.6Q + 0.1Q^2

Shut down point has AVC'(Q) = 0

-1.6 + 0.1*2Q = 0

Q = 1.6/0.2 = 8 units

AVC (minimum) = shut down point = 12 - 1.6*8 + 0.1*8^2 = $5.6

Hence, the shut down price is $5.6